The gig economy has changed the world of work. Discuss

Assume the following national income accounting data in billions of ringgit:
                                                                                                                                  RM billion
transfer payments to persons
400

personal taxes
100

corporate profit tax payments, undistributed profits and valuation adjustments
200

interest paid to businesses
50

contributions to Social Security
350

net personal transfer payments to foreigners
15

personal interest income
150

government purchases of goods and services
100

consumption expenditures
750

gross investment
75

imports of goods and services
20

exports of goods and services
10

Depreciation
250

indirect taxes
100

Calculate gross national product.
         
Calculate net national product.

Calculate national income.

Calculate personal income.

Calculate personal disposable income.

Calculate personal saving.
Answer #1

Gross National Product (GNP)= Gross Domestic Product (GDP) + net factor Income from Abroad (NFIA)

                                                = Consumption Expenditure(C) + Investment expenditure (I) + Governement expenditure

                                                   (G) + Net exports (exports- imports) + NFIA

                                               = 750+75+100+(10-20)+15

                                      GNP= 930

Net National product (NNP) = GNP - Depreciation(D)

                                             = 930-250

                                             = 580

National Income (NI) = NNP- indirect taxes +subsidies

                                  = 580-100+0

                                  =480

Personal Income (PI)= NI- Corporate tax -undistributed profits- net interest payments made by households+ transfer payments

                             PI= 480-200-50+400

                             PI= 630

Personal disposable Income (DI)= PI - personal tax payments - non tax payments

                                                    = 630-100-350

                                                DI= 180

Now,

               DI= savings +expenditure

               savings= DI- 150

                            = 180-150

                            =30